OpenStudy (anonymous):

Find the solution of the following system of linear equations. y = x - 2 y = -2x + 4

OpenStudy (aroub):

By substitution: x-2=-2x+4 2x+x=2+4 3x=6 x=2 y=x-2 y=2-2 y=0

OpenStudy (anonymous):

very good answer Thanks!!!!!